OpenStudy (anonymous):

find the range of the function f(x) = definite integral -4 to x sqrt(16 - t^2) dt.

OpenStudy (anonymous):

[-4, 4] [-4, 0] [0,-4] [0,4pi] [0, 8pi]

OpenStudy (anonymous):

\[I=\int\limits_{-4}^{x}\sqrt{16-t^2} dt\] for range \[\sqrt{16-t^2}\ge 0,16-t^2\ge 0,16\ge t^2\] t^2 <=16 \[\left| t \right|<=4 \] \[\left[ -4,4 \right]\]

OpenStudy (anonymous):

are you sure that isn't the domain?

OpenStudy (anonymous):

no ,it is domain.

OpenStudy (anonymous):

\[I=\int\limits_{-4}^{x}\sqrt{16-t^2}*1~dt=[ \sqrt{16-t^2}* t]({-4 \rightarrow x})-\int\limits_{-4}^{x}\frac{ -2t }{2\sqrt{16-t^2} }*t*dt\] \[=\sqrt{16-x^2}-0-\int\limits_{-4}^{x}\frac{ 16-t^2-16 }{\sqrt{16-t^2} }dt\] \[=\sqrt{16-x^2}-\int\limits_{-4}^{x}\sqrt{16-t^2}dt+16\int\limits_{-4}^{x}\frac{ dt }{ \sqrt{4^2-t^2} }\] \[I+I=\sqrt{16-x^2}+16\sin^{-1} \frac{ t }{ 4 }\left({-4 \rightarrow x} \right)\] \[I=\frac{ \sqrt{16-x^2} }{ 2 }+8\left( \sin^{-1} \frac{ x }{4 }-\sin^{-1} \frac{ -4 }{4 } \right)\]

OpenStudy (anonymous):

i have to go.

OpenStudy (anonymous):

Think about what the integrand represents: \(\sqrt{16-t^2}\). You know \([-4,4]\) is the domain of \(f(x)\). Since \(t\) is just a placeholder, let's call it \(x\) for a moment. Plotting this, you have